a farmer has $2400 to spend to fence two rectangular pastures. the local contractor will build the fence at a cost of $6.25/m what is the largest total area that the farmer can have fenced for that price ?
Answer by rfer(16322)

2400/6.25=384 m
384/4=96
96^2=9216 sq m